# Toxic Bias: Perspective API Misreads German as More Toxic

### v1

@@ 4.2.1. Toxicity of tweets and users
4.2.2. Toxicity of tweets translated in English
To rule out the possibility that the higher toxicity scores assigned to German texts are due to more toxic content, we translate both German and Italian tweets into English and we subsequently use Perspective API to compute toxicity scores for the English texts. We leverage Argos Translate (argotranslate, ) to obtain English translations of our tweets. Figure 4 shows the distribution of toxicity scores for all German tweets (solid line) and their English translation equivalents (dashed line). We observe that the tweets translated in English are much less toxic, with their distribution being similar to that of generic tweets in non-German languages, presented in Figure 1.
Figure 4. Distribution of toxicity scores for German COVID-19 vaccine-related tweets and for their English translation. Median values are: DE = 0.132, EN = 0.012.
Figure 5. Distribution of the differences in toxicity scores for German and Italian tweets compared to their English translation. Median values are DE = +0.085, IT = -0.006.
However, it may be that the machine-translation tends to reduce the toxicity of messages, for instance by refusing to translate certain swearwords or slang. To examine whether this is the case, we also compute the difference in toxicity between a random sample of 10 0001000010\,00010 000 tweets in German compared to their respective English translation; we also translate them from German to Italian and compute their difference in toxicity with the English version. Figure 5 shows the distribution of these differences. We observe that, on average, German messages are 8.5% more toxic (i.e., median difference = +0.085) than the corresponding English messages, as also shown by the skewed orange-colored distribution. Contrarily, the differences between Italian and English-translated messages appear to follow a normal distribution centered around 0 and, in fact, Italian messages only show negligible differences (median difference = 0.01) when compared to their English translations. These results unequivocally demonstrate that the larger toxicity scores for German texts do not depend on the content of such texts, but are rather the consequence of an inherent language bias in Perspective API. To rule out potential biases in the translation that might modify the toxicity, we perform the same experiments by starting from a random sample of 10 0001000010\,00010 000 Italian tweets and translating them to German and English to compute differences. Also in this case, we obtained qualitatively similar results, but omit the figure for the sake of brevity.
@@ 4.2.3. Implications for content moderation
Specifically, we focus on how different toxicity thresholds used for removing tweets or users might affect COVID-19 vaccine conversations in German when compared to their translation in English, using the random sample of 10 0001000010\,00010 000 tweets employed in the previous analysis. In the top panel of Figure 6, we simulate removing all tweets whose toxicity score exceeds a given threshold, and we show the percentage of extra tweets removed when considering German scores compared to English scores. For instance, by adopting the threshold value of $0.7$ recommended by Perspective API (thresholdperspective, ), 10 times more German tweets would be removed, and similarly for the number of German users (as shown in the bottom panel of the figure). The median increase in percentage points is 429.94 for tweets and 409.84 for users, meaning that on average more than four times German tweets and users would be removed compared to English.
Figure 6. Proportion of extra tweets (top) and users (bottom) removed in German compared to English for different choices of toxicity score as a threshold, i.e., tweets/users above a given score would be removed. The median increase in percentage points is 429.94 for tweets and 409.84 for users, meaning that on average more than 4 times German tweets and users would be removed compared to English.
@@ 4.3. Wikipedia texts in German are more toxic than in English
Figure 7. Distribution of toxicity for 12 0001200012\,00012 000 random Wikipedia page summaries in English and German language. Median values are: DE = 0.03, EN = 0.01.
@@ 5. Discussion
Perspective API has emerged as the technological foundation for a broad spectrum of applications, stretching from academic research to practical implementations in content moderation across various online platforms. In the realm of research, scholars and data scientists utilize Perspective API to analyze online discourse, study patterns of hate speech and harassment, and explore the dynamics of digital communication. Beyond the academic sphere, Perspective API plays a pivotal role in content moderation for websites, forums, and large-scale social media platforms. By automatically flagging potentially harmful or abusive comments, it aids moderators in maintaining a constructive and respectful environment, enabling large-scale platforms to manage large volumes of user-generated content that would make manual moderation impractical. Perspective API is often integrated into comment systems and discussion forums, enabling preemptive identification and removal of messages that are classified as toxic. While the widespread application of the API means that the system provides important support for researchers and platforms, it also highlights the potential negative impact of any biases or inaccuracies in its algorithm.
This paper has identified a significant bias within the model when it comes to processing German-language messages: our analysis reveals that the algorithm tends to assign disproportionately higher toxicity scores to text written in German, compared to similar content in other languages. Specifically, our examination of Twitter posts revealed that the same content reported in German was more toxic than its English-translated version. This pattern was further validated by analyzing random texts from Wikipedia summaries in both languages. The distribution of toxicity returned by the API for German appears spiky, rather than following the smooth distribution of other languages – suggesting some errors or artifacts stemming from the underlying machine learning model. Due to the black box and proprietary nature of the model, we unfortunately cannot examine the underlying reasons for these results.
@@
The problems identified in this paper highlight the inherent risks as researchers become increasing dependent on proprietary models for data analysis and AI. The closed and proprietary APIs offered by major tech companies provide researchers with easy access to vast amounts of data and sophisticated analysis methods, which can otherwise be challenging to access. This ready availability is invaluable for studying complex social phenomena, trends, and behaviors in the digital age. However, the growing dependency on closed models also raises critical concerns. The opaque nature of these proprietary systems means that researchers are unable to scrutinize or understand the inner workings of the algorithms, and identify potential biases and limitations. The result is a substantial risk of systemic biases being embedded and perpetuated through their employment. This lack of transparency can furthermore inadvertently influence the direction and conclusions of studies, as researchers must operate within the constraints and methodologies predefined by these tools. While proprietary APIs have opened new avenues for exploring social sciences, they also necessitate a cautious and critical approach to ensure the integrity of research and its reproducibility (pozzobon2023challenges, ) – rather than taking the technology firms by their word.
@@
Ethical considerations. We envision positive ethical implications for this work, particularly concerning future research on online toxicity as well as on the policing of online platforms. Our findings underscore the potential consequences of relying on black-box models and APIs, without a thorough understanding of their language biases. The identification of intrinsic biases within the multilingual Perspective API highlights the risk of inadvertently amplifying or suppressing certain voices in online spaces. The observed higher levels of toxicity assigned to German content compared to other languages raise concerns about the fairness and equity of ongoing content moderation processes, which could potentially lead to disproportionate censorship or scrutiny of specific linguistic communities. This study urges a reevaluation of the ethical dimensions surrounding the deployment of such tools, emphasizing the need for transparency, accountability, and ongoing refinement to ensure that AI-driven content moderation aligns with principles of impartiality and cultural sensitivity, thereby mitigating the inadvertent amplification of biases in online discourse.

### v3

@@ 4.2. German vaccine tweets are more prone to removal
Figure 5. Distribution of toxicity scores for German COVID-19 vaccine-related tweets and for their English translation, from Dataset 2. Histograms are built with 100 equal-width bins. Median values are: DE = 0.132, EN = 0.012.
@@ 4.2.1. Toxicity of tweets and users
4.2.2. Toxicity of tweets translated to English
To rule out the possibility that the higher toxicity scores assigned to German texts are due to more toxic content, we translate both German and Italian tweets into English and we subsequently use Perspective API to compute toxicity scores for the English texts. We leverage Argos Translate [^4] http://www.argosopentech.com/ to obtain English translations of our tweets. Figure 5 shows the distribution of toxicity scores for all German tweets (solid line) and their English translation equivalents (dashed line). We observe that the tweets translated in English are much less toxic, with their distribution being similar to that of generic tweets in non-German languages, presented in Figure 1.
Figure 6. Distribution of the differences in toxicity scores for German and Italian tweets compared to their English translation, using a random sample of 10 0001000010\,00010 000 tweets from Dataset 2. Histograms are built with 100 equal-width bins. Median values are DE = +0.085, IT = -0.006.
However, it may be that the machine-translation will tend to reduce the toxicity of messages, for instance by avoiding to translate certain swearwords or slang. To examine whether this is the case, we also compute the difference in toxicity between a random sample of 10 0001000010\,00010 000 tweets in German compared to their respective English and Italian translation. Figure 6 shows the distribution of these differences. We observe that, on average, German messages are 8.5% more toxic (i.e., median difference = +0.085) than the corresponding English messages, as also shown by the skewed orange-colored distribution. Contrarily, the differences between Italian and English-translated messages appear to follow a normal distribution centered around 0 and, in fact, Italian messages only show negligible differences (median difference = 0.01) when compared to their English translations. These results unequivocally demonstrate that the larger toxicity scores for German texts do not depend on the content of such texts, but are rather the consequence of an inherent language bias in Perspective API. To rule out potential biases in the translation that might modify the toxicity, we perform the same experiments by starting from a random sample of 10 0001000010\,00010 000 Italian tweets and translating them to German and English to compute differences. Also in this case, we obtained similar results, but omit the figure for the sake of brevity.
4.3. Correlation between toxicity and difference with English translation
[…]
Figure 7. Distributions of difference in toxicity between German tweets from Dataset 2 and their English translation, divided by 20 bins of toxicity computed on the German version of each tweet. Outliers are not shown and triangles indicate the mean of the distribution.
@@ 4.3.1. Implications for content moderation
Specifically, we focus on how different toxicity thresholds used for removing tweets or users might affect COVID-19 vaccine conversations in German when compared to their translation in English, using the random sample of 10 0001000010\,00010 000 tweets employed in the previous analysis. In the top panel of Figure 8, we simulate removing all tweets whose toxicity score exceeds a given threshold, and we show the percentage of extra tweets removed when considering German scores compared to English scores. For instance, by adopting the threshold value of $0.7$ recommended by Perspective API [^5] https://developers.perspectiveapi.com/s/about-the-api-score?language=en_US, 10 times more German tweets would be removed, and similarly for the number of German users (as shown in the bottom panel of the figure). The median increase in percentage points is +429.94 for tweets and +409.84 for users, meaning that on average more than four times German tweets and users would be removed compared to English.
Figure 8. Proportion of extra tweets (top) and users (bottom) removed in German compared to English for different choices of toxicity score as a threshold, i.e., tweets/users above a given score would be removed. The median increase in percentage points is +429.94 for tweets and +409.84 for users, meaning that on average more than 4 times German tweets and users would be removed compared to English.
@@ 5. Discussion
5.1. Contributions
Perspective API has emerged as the technological foundation for a broad spectrum of applications, stretching from academic research to practical implementations in content moderation across various online platforms. In the realm of research, scholars and data scientists utilize Perspective API to analyze online discourse, study patterns of hate speech and harassment, and explore the dynamics of digital communication. Beyond the academic sphere, Perspective API plays a pivotal role in content moderation for websites, forums, and large-scale social media platforms. By automatically flagging potentially harmful or abusive comments, it aids moderators in maintaining a constructive and respectful environment, enabling large-scale platforms to manage large volumes of user-generated content that would make manual moderation impractical. Perspective API is often integrated into comment systems and discussion forums, enabling preemptive identification and removal of messages that are classified as toxic. For these reasons, automated tools such as Perspective API apparently contribute to overcoming the challenges in content moderation and to respecting the regulatory obligations that large online platforms must abide by (trujillo2023dsa, ). However, while the widespread application of such tools means that these systems provide important support for researchers and platforms alike, it also highlights the potential negative impact of any biases or inaccuracies in the data and algorithms.
Figure 9. Distribution of toxicity for over 10 0001000010\,00010 000 random Wikipedia page summaries in Arabic (ar), Chinese (zh), English (en), German (de) and Japanese (ja). Histograms are built with 100 equal-width bins. We limit the x-axis to 0.3 for clarity reasons, as the distribution is highly skewed towards 0.
This paper identified a significant bias within the Perspective API model when it comes to processing German-language messages: our analysis reveals that the algorithm tends to assign disproportionately higher toxicity scores to texts written in German, compared to similar content in other languages. Specifically, our examination of Twitter posts revealed that the same content reported in German was more toxic than its English-translated version. This pattern was further validated by analyzing random texts from Wikipedia summaries in both languages. The distribution of toxicity returned by the API for German appears spiky, rather than following the smooth distribution of other languages – suggesting some errors or artifacts stemming from the underlying machine learning model. Due to the black box and proprietary nature of the model, we unfortunately cannot examine the underlying reasons for these results, which calls for additional research efforts. However, an exploratory analysis showed that no specific characters and words are associated with such spiky behaviour, and that no abnormal patterns are observed for other non-European languages such as Arabic, Chinese and Japanese.
@@ 5.1. Contributions
The problems identified in this paper highlight the inherent risks as researchers become increasingly dependent on proprietary models for data analysis and AI. The closed and proprietary APIs offered by major tech companies provide researchers with easy access to vast amounts of data and sophisticated analysis methods, which can otherwise be challenging to access. This ready availability is invaluable for studying complex social phenomena, trends, and behaviors in the digital age. However, the growing dependency on closed models also raises critical concerns. The opaque nature of these proprietary systems means that researchers are unable to scrutinize or understand the inner workings of the algorithms, and identify potential biases and limitations. The result is a substantial risk of systemic biases being embedded and perpetuated through their employment. This lack of transparency can furthermore inadvertently influence the direction and conclusions of studies, as researchers must operate within the constraints and methodologies predefined by these tools. While proprietary APIs have opened new avenues for exploring social sciences, they also necessitate a cautious and critical approach to ensure the integrity of research and its reproducibility (pozzobon2023challenges, ) – rather than taking the technology firms by their word.
5.2. Limitations
There are some limitations in our work. We based our investigation on data from Twitter, which is not the most widely adopted social platform in many countries, and therefore not fully representative of the population. However, our robustness analyses with Wikipedia data suggest that the highlighted issues do not depend on user demographics or the specific context/topic of conversation. Besides, we only analyzed the Toxicity attribute of the API, and leave an investigation of other attributes (such as Insult or Profanity) for future research. Lastly, We did not evaluate whether this bias manifests in other LLM-based models for toxicity detection.
5.3. Ethical considerations
@@
We envision positive ethical implications for this work, particularly concerning future research on online toxicity as well as on the policy design of online platforms. Our findings underscore the potential consequences of relying on black-box models and APIs, without a thorough understanding of their language biases. The identification of intrinsic biases within the multilingual Perspective API highlights the risk of inadvertently amplifying or suppressing certain voices in online spaces. The observed higher levels of toxicity assigned to German content compared to other languages raise concerns about the fairness and equity of ongoing content moderation processes, which could potentially lead to disproportionate censorship or scrutiny of specific linguistic communities. This study urges a reevaluation of the ethical dimensions surrounding the deployment of such tools, emphasizing the need for transparency, accountability, and ongoing refinement to ensure that AI-driven content moderation aligns with principles of impartiality and cultural sensitivity, thereby mitigating the inadvertent amplification of biases in online discourse.
5.4. Data availability
To ensure reproducibility, we will provide access to the code and datasets employed in our analysis in a public repository after the revision process. For what concerns Dataset 1 and Dataset 2, to comply with Twitter’s Terms of Service, we can only provide tweet IDs, which can be employed (hydrated) to retrieve the data by means of Twitter API. We acknowledge that access to the GET /2/tweets endpoint for researchers is not as easy as in the past, and we encourage interested researchers to reach out to the authors in case they encounter difficulties in replicating the analyses. For what concerns Dataset 3, we will provide the full data, as it is already publicly accessible. We notice that we do not attempt to de-identify or disclose any personal or sensitive data.
6. Appendix
Here we provide supplementary figures and results referenced and commented in the main text.
Figure 10 provides the distribution of toxicity scores for tweets shared in German-speaking countries from Dataset 1 separating them by the number of non-ASCII characters they contain.
Figure 11 provides the classification performance, computed with a 5-fold cross-validation, of three off-the-shelf classifiers in the binary task of predicting whether a tweet belongs to the Abnormal class described in the main text.
Figure 12 provides the distributions of toxicity scores for tweets containing (or not) one of the top 10 words differentiating the Abnormal class (in terms of Z-score). They are translated into English as follows:
  • nicht: not
  • die: the
  • sie: they/she
  • man: one
  • ist: is
  • zu: to
  • sich: oneself
  • wenn: when
  • nur: only
  • dass: that
Figure 10. Distributions of toxicity scores for tweets from German-speaking countries in Dataset 1 based on the number of non-ASCII characters they contain.
Figure 11. Performance evaluation of three classifiers in the binary task of predicting whether a tweet belongs to the class of Abnormal toxicity scores described in the main text. We show the mean value of each metric in a 5-fold cross-validation.
Figure 12. Distributions of toxicity scores for tweets from German-speaking countries in Dataset 1 that contain (or not) one of the top-10 words differentiating the Abnormal class.